OpenStudy (anonymous):

Please help I don't get this What is the slope of the line that passes through the points (–2, 1) and (1, 6)? A. -5 B. 3/5 C. -7 D. 5/3

OpenStudy (kitten_is_back):

which one do you think is you answer love?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

maybe b

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

how? can u please show me your attempt?

OpenStudy (anonymous):

yeah @hartnn -2-1/1-6 = -3/5

hartnn (hartnn):

oh you interchanged x and y values numerator is difference of y values 6-1 = .. ? denominator is difference of x values 1 - (-2) = .. ?

OpenStudy (anonymous):

6-1-5 1-(-2)=3 @hartnn

hartnn (hartnn):

good! so m=5/3 :)

OpenStudy (anonymous):

thank you soo much @hartnn

hartnn (hartnn):

welcome ^_^